# A Fast, Analytic Empirical Model of the Gaia Data Release 3 Astrometric Orbit Catalog Selection Function

### v1

@@ 4.3 Lack of sensitivity at $P\approx 1$ year and lack of solutions with $P>550$ days
We test this hypothesis by cross-matching the nss_acceleration_astro  catalog to external binary catalogs to empirically estimate how many orbital solutions are lost to acceleration solutions in the model cascade (Appendix B). After accounting for the number of solutions at a given orbital period, we confirm that the number of acceleration solutions increases at longer periods, and the number of orbital solutions decreases at longer periods. This is also confirmed by the modeling by El-Badry et al. (2024) (their Figure 11).
Modeling the dearth of solutions at $P_{orb}\approx 1$ year and the decrease in number of long $P_{orb}$ solutions require detailed knowledge of the spacecraft scanning law and individual epoch data. In the spirit of keeping calculations fast and analytic in this work, we construct a probabilistic function to empirically account for these effects. For $P_{orb}<250$ days, all solutions are kept; we assume that the $\texttt{ruwe}>1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters fully capture the astrometric cascade. [^9] This is not entirely accurate; examination of the period distribution show that solutions are clearly lost around 190 days, which is 3×3$\times$3 × Gaia’s precession period. However, the number of solutions is so small we ignore this. We then assume the remaining discrepancy between the Gaia NSS astrometric orbits catalog and the forward model period distributions are due to the scanning law, finite Gaia DR3 temporal baseline, and loss of orbital solutions to acceleration solutions. We empirically determine this function by dividing the number of NSS orbital solutions by the number of solutions in the forward model after application of the ruwe  and $\varpi/\sigma_{\varpi}$ filters. We normalize the distribution assuming no solutions are lost at $P_{orb}=250$ days.
Figure 5 shows the probabilistic function as a function of orbital period. We do not attempt to quantitatively understand the particular form of this function as this requires epoch-level modeling, which is outside the scope of this work. We also note that this empirical function is a function of $P_{orb}$ alone, which again is not completely accurate; for example, the 1 year orbit degeneracy can be more easily broken by highly eccentric orbits than circular orbits.
@@ 5.1.1 Number of solutions
Table 1 summarizes the effect of the filters on the number of orbital solutions. The forward model begins with 150 million binaries within 2 kpc of the Sun at the present day. Even before applying the filters discussed in the previous section, many binaries are simply not detectable because they are too faint ($G\lesssim 21$) to be detected by Gaia, or their orbital periods are very long ($P_{orb}\gtrsim$ 30 years). In order to generate a more realistic estimate on the number of potentially detectable Gaia binaries, we calculate the number of binaries with $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which represents the range of orbital periods and magnitudes spanned in the NSS astrometric orbits catalog; 13 million binaries satisfy these criteria. Ultimately, after applying all the filtering cuts, we find $3.5\times 10^{5}$ binaries will be included in the NSS catalog.
@@
This is double the number of sources actually included in the NSS astrometric orbits catalog (168,065). There are several possible explanations for this overestimate. One possibility is that the initial binary population produced by the binary population synthesis was incorrect in some way, e.g., the absolute number of systems was too high, the orbital period distribution of the binaries were incorrect and produced too many systems with $P_{orb}\approx 100-1000$ days, or the extinction was underestimated. Another is that the filtering steps do not accurately represent the model cascade, e.g., estimating the uncertainties $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ using the NSS astrometric catalog itself biases their values low. Nonetheless, the forward model is able to reproduce the total number of sources within a factor of two, which will still provide useful constraints in population studies.
Table 1: NSS filters and resultant number of predicted binaries
Filter No. of binaries (%)
None (all binaries) $1.5~{}\times~{}10^{8}$ (100%)
$15<P_{orb}<10^{4}$ day, $18.54<G<$ 18.5 mag $1.3~{}\times~{}10^{7}$ (8.8%)
ruwe$>1.4$ $6.3~{}\times~{}10^{6}$ (4.2%)
$\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ $1.2~{}\times~{}10^{7}$ (8.2%)
$a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{ day}})$ $8.6~{}\times~{}10^{6}$ (5.8%)
ruwe$>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ $5.6~{}\times~{}10^{6}$ (3.7%)
ruwe$>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$, $P_{orb}\approx$ 1 yr, $P_{orb}>550$ day $3.5~{}\times~{}10^{5}$ (0.2%)
Note. — The first line, where Filter is “None”, denotes the number of binaries within 2 kpc of the Sun. The filter on $15<P_{orb}<10^{4}$ day, $18.54<G<$ 18.5 mag is not an actual filter criterion for the NSS astrometric orbits catalog, but is a more representative number of the possible number of orbital systems that could be detected by Gaia as compared to all binaries within 2 kpc, which has many systems with orbital periods too long or magnitudes too faint to be detected.
@@ 5.1.2 Distribution of binary and orbital parameters
Figure 6: Effect of filtering on $\texttt{ruwe}>1.4$ (ruwe), $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ ($\varpi/\sigma_{\varpi}$), $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ ($a_{0}/\sigma_{a_{0}})$, and $P_{orb}\approx 1$ yr and $P_{orb}>550$ day ($P$) on the period distribution in the forward modeling.
Figure 7: 1-D and 2-D PDFs in the period $P_{orb}$, parallax $\varpi$, eccentricity $e$, photocenter orbit semimajor axis $a_{0}$, ruwe, apparent magnitude $G$, and parallax over parallax error $\varpi/\sigma_{\varpi}$ of the forward model (red) as compared to the actual NSS astrometric orbits catalog (black). In general, the forward modeling does a good to reasonable job in matching the actual NSS distributions including the correlations, with the exception of the eccentricity distribution.
@@
Figure 8: Effect of filtering on the distribution of orbital period $P_{orb}$, parallax $\varpi$, eccentricity $e$, ruwe, magnitude $G$, primary and secondary mass $M_{1},M_{2}$, luminosity ratio $l_{G}$, mass ratio $q$, and photocenter scale $\delta_{ql}$. The dotted line shows the birth distribution from Moe & Di Stefano (2017). The black outline shows the distribution of systems that have orbital periods between $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia. The filter on long-period orbits preferentially removes larger photocenters and large ruwe. The magnitude cut preferentially removes distant and low-mass binaries. It causes the primary mass distribution to peak roughly around $0.8M_{\odot}$ (when binned logarithmically). The gray outline shows the systems that satisfy $\texttt{ruwe}>1.4$. Finally, the red outline shows the systems that satisfy all the filtering criteria applied ($\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,\sqrt{P_{orb}/\textrm{days}})$, $P_{orb}\approx 1$ year, $P_{orb}>550$ day).
Figure 6 summarizes the effect of the filtering steps on the period distribution. The $\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, and $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ filters all preferentially remove shorter period systems. The ruwe  filter removes more systems overall, while $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$ filters remove shorter period solutions more strongly.
@@
Figure 7 is a corner plot comparing the orbital period, parallax, eccentricity, photocenter orbit semimajor axis, ruwe, apparent magnitude, and parallax over parallax error of the mock forward model to the actual NSS astrometric orbit catalog. The 2-D correlations are broadly preserved, and the 1-D distributions are also generally reasonable matches. The deficit of solutions around $P_{orb}\approx 1$ year and the decline in number of solutions longer than $P_{orb}>550$ days in the forward modeled period distribution were tuned to match the NSS catalog, so they are unsurprisingly a reasonable match. The parallax, magnitude, and parallax over parallax error distributions are also a very good match. The photocenter orbit semimajor axis and ruwe  distributions are also reasonable matches, although the models tend to result in distributions slightly more sharply peaked than the NSS catalog. The eccentricity distribution between the model and the true catalog are by far the most discrepant, with the Gaia NSS distribution peaking around $e\approx 0.2$, while the forward modeled distribution peaks at a much higher eccentricity of $e\approx 0.75$. We discuss possible reasons for this in §6.1. Although the parallax error and photocenter orbit semimajor axis errors were interpolated directly using the NSS catalog itself, the filters and distributions involving $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ were not tuned post-facto like the period distribution.
Figure 8 compares the 1-D distribution of binary parameters in the initial birth population with that obtained from a simple filter on observability, which we define to be $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia. It also shows the distribution of parameters after application of the $\texttt{ruwe}>1.4$ filter and all NSS filters. As discussed previously, the ruwe  filter trims short orbital period, small $a_{0}$, distant, and faint systems. Including the rest of the filters does not significantly change the 1-D distribution of the parameters as compared to the $\texttt{ruwe}>1.4$ filter, except the orbital period distribution.

### v2

@@ 4.3 Removal of acceleration solutions candidates
As an independent test of whether or not longer period orbital solutions are preferentially assigned acceleration solutions, we cross-match the nss_acceleration_astro  catalog to external binary catalogs to empirically estimate how many orbital solutions are lost to acceleration solutions in the model cascade (Appendix D). After accounting for the number of solutions at a given orbital period, we confirm that the fraction of acceleration solutions increases at longer periods, and the fraction of orbital solutions decreases at longer periods.
4.4 Orbital model acceptance criteria: parallax and photocenter semimajor axis error
Figure 7: Parallax errors in the gaia_source (left) and NSS (right) catalogs. The analytic prediction (Equation 7) for the parallax errors and the true catalog errors are tightly correlated in the case of gaia_source (top panel); there is a tail of underestimated errors that we attribute to non-single sources being erroneously fit with single-star models (bottom panel). On the other hand, the analytic prediction and the true catalog errors are highly discrepant in the NSS catalog; the correlation is weak (top panel) and the analytic estimates systematically underestimate the errors (bottom panel). This is due to erroneous attribution of errors due to correlations between the parallax and orbit model parameters such as orbital period. This poor agreement necessitates an empirical correction to $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ in order to provide accurate uncertainties for NSS sources (Figure 8).
In order to model the selection criteria of the orbital solutions, we need an estimate of the orbit significance $s_{orb}=a_{0}/\sigma_{a_{0}}$ and the parallax over error $\varpi/\sigma_{\varpi}$. The parameters $a_{0}$ and $\varpi$ are directly calculated from the binary source’s properties; we thus need a way to estimate their uncertainties.
4.4.1 Analytic $\sigma_{\varpi},\sigma_{a_{0}}$ prediction
The parallax error $\sigma_{\varpi}$ and photocenter orbit semimajor axis error $\sigma_{a_{0}}$ are expected to be proportional to the individual measurement errors, and decrease with the square root of the number of observations: $\sigma_{AL}(G)/\sqrt{N_{vis}}$. Using a random subset of the single-star fits in the gaia_source catalog, we find that the constant of proportionality is 0.48 for the parallax errors (Figure 7). Because we cannot measure the constant of proportionality for $\sigma_{a_{0}}$ using the gaia_source catalog since it consists of single-star fits, we indirectly infer it by inspecting the distribution of $\log_{10}(\sigma_{\varpi}/\sigma_{a_{0}})$ in the NSS catalog. It is symmetric with its peak at $\sigma_{\varpi}/\sigma_{a_{0}}\approx 1$, so we assume that the constant of proportionality for $\sigma_{a_{0}}$ is the same as for $\sigma_{\varpi}$. We thus adopt for our analytic error model
$\sigma_{\varpi},\sigma_{a_{0}}=\frac{0.48\sigma_{AL}(G)}{\sqrt{N_{vis}}}.$ (7)
Having the constant of proportionality exactly correct is not crucial, since we will be empirically rescaling it, as discussed in the next section.
4.4.2 Empirical correction to analytic $\sigma_{\varpi},\sigma_{a_{0}}$ prediction
In practice, Equation 7 will underestimate the true $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ for sources in the NSS astrometric orbits catalog (Figure 7). Analogous to the discussion in §4.2.2, this underestimation is due to the erroneous correlations between the parallax, position, proper motion and orbital parameters when performing the model fit. This erroneous error cannot be derived analytically, as it depends on the modeling of individual epoch observations. Thus, we construct an empirical $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ correction using the simulations of El-Badry et al. (2024), analogous to the empirical ruwe correction in §4.2.2.
We analyze $3.3\times 10^{5}$ sources from the simulated cascade of El-Badry et al. (2024) that are fit with an orbital solution and that have reliably measured periods and eccentricities (i.e., the inferred $P_{orb}$ and $e$ are consistent within $3\sigma$ of the true $P_{orb}$ and $e$). Although the simulated cascade of El-Badry et al. (2024) fits $2.2\times 10^{6}$ orbital solution candidates, we only model the 15% of these sources with reliable fits as the remainder are spurious solutions that are not going to have accurate error estimates. Because El-Badry et al. (2024) are modeling the full Gaia DR3 cascade, they generate spurious solutions which are removed by the $F_{2}$ cut. These spurious solutions are not present in our analytic model, as they arise purely from the epoch modeling.
We bin the reliably fit sources by their orbital periods and eccentricities. Within each bin, we fit a Gaussian to the distribution of $\log_{10}(\sigma_{\varpi,fit}/\sigma_{\varpi,analytic})$, where $\sigma_{\varpi,fit}$ is the parallax error as measured when modeling epoch astrometry in the catalog of El-Badry et al. (2024), and $\sigma_{\varpi,analytic}$ is the parallax error predicted from Equation 7. We then interpolate over this grid of values to sample the means and standard deviations, which are then sampled to assign a correction to $\sigma_{\varpi,analytic}$. We perform the same procedure on $\sigma_{a_{0},fit}$ and $\sigma_{a_{0},analytic}$ to determine a correction for the photocenter orbit semimajor axis $\sigma_{a_{0},analytic}$.
The empirical correction is shown in Figure 8. The correction for $\sigma_{\varpi}$ is primarily a function of the orbital period. Equation 7 underpredicts $\sigma_{\varpi}$ by a factor of $10^{0.7}\approx 5$ for $P_{orb}\approx 1$ year, due to the degeneracy between $P_{orb}=1$ year orbits and parallax. The correction for $\sigma_{a_{0}}$ is both a function of eccentricity and orbital period. Solutions with either long orbital periods $P_{orb}\gtrsim 1000$ day or high eccentricities $e\gtrsim 0.6$ have $\sigma_{a_{0}}$ corrections around two orders of magnitude larger than solutions with lower eccentricity and shorter orbital period, excluding orbits with $P_{orb}\approx 1$ year. This demonstrates that the significance of the orbital solution is a function of the properties of the orbit itself.
We emphasize that these empirical corrections to $\sigma_{a_{0}}$ and $\sigma_{\varpi}$ are only for solutions that have not received an acceleration solution in the NSS cascade, as they are derived only using sources that are fit with an orbital solution. We also emphasize that the corrections are specific to the observing strategy of Gaia DR3, and that they cannot be generally applied to observations that have a different temporal sampling, including Gaia DR4.
Figure 8: Difference in predicted parallax error $\sigma_{\varpi}$ (left) and photocenter orbit semimajor axis $\sigma_{a_{0}}$ (right), as a function of orbital period, eccentricity, and photocenter orbit semimajor axis for a subset of bright, well-measured sources. The required rescaling of $\sigma_{\varpi}$ is primarily driven by orbital period, with it being significantly under-predicted at $P_{orb}\approx 1$ year, as well as being somewhat under-predicted at longer orbital periods ($P_{orb}\gtrsim 550$ days). The rescaling of $\sigma_{a_{0}}$ is more complex, with dependence on both orbital period and eccentricity. In addition to being under-predicted at $P_{orb}\approx 1$ year and $P_{orb}\gtrsim 1000$ day, it is also significantly underpredicted at high eccentricities ($e\gtrsim 0.6$).
@@ 5.1.1 Number of solutions
Table 1 summarizes the effect of the filters on the number of orbital solutions. The forward model begins with 150 million binaries within 2 kpc of the Sun at the present day. Even before applying the filters discussed in the previous section, many binaries are simply not detectable because they are too faint ($G\gtrsim 21$) to be detected by Gaia, or their orbital periods are very long ($P_{orb}\gtrsim$ 30 years). In order to generate a more realistic estimate on the number of potentially detectable Gaia binaries, we calculate the number of binaries with $P_{orb}<1500$ days and $G<19$ mag, which represents the range of orbital periods and magnitudes spanned in the NSS astrometric orbits catalog; 12 million binaries satisfy these criteria. Ultimately, after applying all the filtering cuts, we find $2.6\times 10^{5}$ binaries will be included in the NSS catalog.
@@
This is about 50% more than the number of sources actually in the NSS astrometric orbits catalog (168,065). There are several possible explanations for this overestimate. For example, the initial binary population produced by the binary population synthesis was incorrect in some way, e.g., the rescaling factor in §3 based on the GCNS was overestimated due to an overdensity of stars in the solar neighborhood, the orbital period distribution of the binaries was incorrect and produced too many systems with $P_{orb}\approx 100-1000$ days, the extinction was underestimated, or some binaries were disrupted by stellar encounters or cluster dispersal. Nonetheless, the forward model is able to reproduce the total number of sources within a factor of two, which will still provide useful constraints in population studies.
Table 1: Selection effect of NSS cascade filters
Filter No. of binaries (%) Selection effect
None (all binaries) $1.5~{}\times~{}10^{8}$ (100%) N/A
NSS potentially detectable ($P_{orb}<1500$ day, $G<19$ mag) $1.2~{}\times~{}10^{7}$ (8%) Preferential selection of primaries with $M_{1}\approx 1M_{\odot}$
NSS catalog candidates: ruwe$>1.4$ $1.3\times 10^{6}$ (0.9%) Preferentially selects $P_{orb}\approx 10^{3\pm 1}$ days
Orbital solution candidates (i.e., not selected to acceleration solution candidates): $s_{7,9}<12$ or $\omega/\sigma_{\omega}>2.1s_{7,9}^{1.05}$ $7.4\times 10^{5}$ (0.5%) Preferentially removes $P_{orb}\gtrsim 500$ days, increasing from 0% at $P_{orb}\approx 500$ days up to 85% at $P_{orb}\approx 2000$ days
Orbital solutions ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}}$) $2.6\times 10^{5}$ (0.2%) Preferentially removes 1) high eccentricity $e\gtrsim 0.6$ solutions; 2) short-period $P_{orb}\lesssim 500$ days; 3) $P_{orb}\approx 1$ year solutions
Note. — Steps in the NSS astrometric processing cascade, the resulting number of binaries, and a summary of the selection effect introduced by the filter. Note that “NSS potentially detectable” is not a filter described in Halbwachs et al. (2023). It is included as a proxy for the fraction of the total binary population that could realistically receive an astrometric solution; most binaries are too faint to be detected by Gaia or have orbital periods significantly longer than the $\approx 3$ year time baseline of DR3.
@@ 5.1.2 Distribution of binary and orbital parameters
Figure 9: Effect of filtering on the distribution of orbital period $P_{orb}$, parallax $\varpi$, eccentricity $e$, ruwe, magnitude $G$, primary and secondary mass $M_{1},M_{2}$, luminosity ratio $l_{G}$, mass ratio $q$, and photocenter scale $\delta_{ql}$. The dashed gray line shows the birth distribution from Moe & Di Stefano (2017). The dotted gray line shows the distribution of systems that have orbital periods $P_{orb}<1500$ days and $G<19$ mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia. The magnitude cut preferentially removes distant and low-mass binaries. It causes the primary mass distribution to peak roughly around $0.8M_{\odot}$ (when binned logarithmically). The solid black line shows sources that are NSS model candidates ($\texttt{ruwe}>1.4$), the blue line shows sources that are orbital model candidates (i.e., NSS model candidates that are not candidates for an acceleration solution), and the red line shows the sources that are in the final astrometric orbits catalog. Zoom-ins of the $P_{orb}$ and $e$ panels are shown in Figure 10.
Figure 10: Effect of the different filters in the NSS processing cascade on the distribution of orbital periods (left) and eccentricity (right). Each line shows a different filter: $\texttt{ruwe}>1.4$ (black dotted), orbit candidates (gray dashed), and the final orbit catalog (red solid). The filters in the acceleration cascade preferentially remove long period orbital solutions. In the orbital cascade, the filter on $\varpi/\sigma_{\varpi}>20,000\textrm{ d}/P_{orb}$ preferentially removes short period and 1-year period solutions, while the filter on $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{days}})$ also preferentially removes eccentric solutions.
@@
Figure 2 is a corner plot comparing the orbital period, parallax, eccentricity, photocenter orbit semimajor axis, ruwe, apparent magnitude, and parallax over parallax error of the mock forward model to the actual NSS astrometric orbit catalog. The 2-D correlations are preserved, and the 1-D distributions are also generally reasonable matches. The orbital period distribution is a good match between the forward model and actual NSS catalog. The distribution of parallaxes in the forward model are systematically offset larger than the distribution of NSS parallaxes by about 0.1 dex; otherwise, the agreement is also good. The distribution of $a_{0}$ and ruwe are excellent matches. The forward model apparent magnitude distribution is somewhat less peaked and has slightly more bright sources than the NSS distribution, but the match is otherwise good. The distribution of parallax over error in the forward model are systematically offset larger than the distribution of NSS parallaxes by about 0.3 dex; part of this may be related to the parallaxes being high by 0.1 dex. The eccentricity distribution of the model and the true catalog are the most discrepant parameter, with the Gaia NSS distribution peaking around $e\approx 0.2$, while the forward modeled distribution is flatter and has a sharp drop-off in the number of solutions around $e\approx 0.6$; we discuss possible reasons for this in §6.1.
5.2 Cascade selection effects
Next, we explore the selection effects of each part of the cascade and how they produce the resultant distributions in Figure 2, in particular for the orbital period and eccentricity distributions.
Figure 9 compares the 1-D distribution of binary parameters in the initial birth population (dashed gray line) with that obtained from a simple filter on systems that could be astrometrically characterized by Gaia (dotted gray line), defined to be $P_{orb}<1500$ days and $G<19$ mag. It also shows the distribution of parameters after application of the $\texttt{ruwe}>1.4$ filter (black line), removal of acceleration solutions (blue line), and the final NSS astrometric orbit catalog (red line). Table 1 summarizes the effects of each filter on the population of sources. We now discuss in detail the impact of each step in the NSS astrometric cascade.
5.2.1 Input catalog
The initial selection of NSS astrometric candidates primarily introduces a selection effect on the orbital period of the solutions. The $\texttt{ruwe}>1.4$ filter preferentially selects sources with $P_{orb}\approx 1000$ days and the orbital period distribution is roughly symmetric in $\log_{10}(P_{orb})$ (Figure 9). As previously mentioned, our ruwe model is only accurate for $P_{orb}<1500$ days, so we validated that the decrease in number of solutions held out to $P_{orb}=10,000$ days using the catalog of El-Badry et al. (2024). Orbits with $P_{orb}\approx 10^{3\pm 1}$ days are the sweet spot in the ruwe criteria: those with $P_{orb}<100$ days will be too small to have a significant photocenter wobble given the measurement uncertainties; those with $P_{orb}>10,000$ days will lack significant orbital coverage during Gaia DR3 and can be approximately modeled by a single-star fit, given the measurement uncertainties. An orbit of 1000 days maximizes the size of the photocenter wobble and orbital coverage during the survey.
Next, we consider the effects on the eccentricity distribution (Figure 9). Short period orbits are preferentially removed by the ruwe  filter, and most short period orbits $P_{orb}\lesssim 30$ days are circular (Moe & Di Stefano, 2017). Although ruwe  also directly removes high eccentricity orbits (Equation 6), it is a weak effect.
Finally, the ruwe filter also preferentially selects nearby sources with large photocenters and photocenter scales (Figure 9).
5.2.2 Acceleration models
Next, we analyze the effect of the acceleration models. Recall that sources that have an accepted acceleration model during the main NSS processing are not able to be orbital solution candidates.
The acceleration portion of the NSS cascade preferentially removes solutions with $P_{orb}>500$ days (Figure 10). Between $P_{orb}\approx 500-2000$ days, the fraction of NSS astrometric candidates that have accepted acceleration solutions increases from 0 to about 85%. Because long $P_{orb}$ solutions are more likely to only have partial orbital period coverage, the acceleration models are more likely to be accepted; orbits with full coverage are more likely to be poor fits to the acceleration models. The acceleration portion of the NSS cascade explains why the peak of the orbital period distribution is at such a short $P_{orb}\approx 500$ days. In the absence of the acceleration cascade, one would expect high sensitivity to $P_{orb}\approx 700-900$ day sources given the $\approx 1000$ day time baseline of Gaia DR3.
The removal of acceleration solutions does not substantially change the eccentricity distribution (Figure 10).
5.2.3 Orbital model selection criteria
Finally, the last step is selecting the orbital model candidates that will actually be accepted into the final catalog. The orbital model significance filter $a_{0}/\sigma_{a_{0}}>5$ has a very strong effect on the eccentricity distribution, preferentially removing eccentric solutions $e\gtrsim 0.6$. It also preferentially removes long-period solutions $P_{orb}\gtrsim 1000$ days. This is because at a fixed source magnitude and number of visibility periods, solutions which do not have good orbital coverage will have larger uncertainties on $a_{0}$. Similarly, eccentric solutions will also have larger uncertainties on $a_{0}$. The $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{d}}$ filter also preferentially removes eccentric systems, although it preferentially removes short period systems as well because of the $1/\sqrt{P_{orb}}$ dependence.
The parallax significance relative to orbital period filter $\varpi/\sigma_{\varpi}>20,000\textrm{d}/P_{orb}$ does not bias against the eccentricity. It preferentially removes short period solutions due to the $1/P_{orb}$ dependence. It also preferentially removes solutions with $P_{orb}\approx 1$ year because these sources have a very low $\varpi/\sigma_{\varpi}$, due to the large errors from the degeneracy between parallax and orbital motion. Although this effect can be seen in the previous two filtering steps, it is much stronger in this step.